### v1

# UAV-UGV Cooperative Trajectory Optimization and Task Allocation for Medical Rescue Tasks in Post-Disaster Environments

Abstract
In post-disaster scenarios, rapid and efficient delivery of medical resources is critical and challenging due to severe damage to infrastructure. To provide an optimized solution, we propose a cooperative trajectory optimization and task allocation framework leveraging unmanned aerial vehicles (UAVs) and unmanned ground vehicles (UGVs). This study integrates a Genetic Algorithm (GA) for efficient task allocation among multiple UAVs and UGVs, and employs an informed-RRT* (Rapidly-exploring Random Tree Star) algorithm for collision-free trajectory generation. Further optimization of task sequencing and path efficiency is conducted using Covariance Matrix Adaptation Evolution Strategy (CMA-ES). Simulation experiments conducted in a realistic post-disaster environment demonstrate that our proposed approach significantly improves the overall efficiency of medical rescue operations compared to traditional strategies, showing substantial reductions in total mission completion time and traveled distance. Additionally, the cooperative utilization of UAVs and UGVs effectively balances their complementary advantages, highlighting the system’s scalability and practicality for real-world deployment.
Index Terms:
Post-Disaster Rescue; Collaboration of UAV and UGV; Task Allocation; Trajectory Optimization; Genetic Algorithm; Informed-RRT*; CMA-ES.
I Introduction
Earthquakes, floods, hurricanes, and other catastrophic events often destroy roads, bridges, and communication networks, severely limiting traditional emergency response capabilities. Consequently, ensuring timely medical assistance becomes an urgent priority, directly influencing survival rates and recovery outcomes. Gür et al.[1] assess the emergency response to the 2023 Kahramanmaras earthquakes (7.7 and 7.6 magnitude), noting that severe infrastructure damage impeded rescue efficiency and medical aid, adversely affecting survival rates and recovery, emphasizing the need for better preparedness and resilient infrastructure.
UAVs and UGVs have emerged as promising solutions due to their flexibility, rapid deployment, and ability to navigate environments that are hazardous or inaccessible to human responders. UAVs offer advantages such as aerial reconnaissance, swift delivery, and the ability to bypass ground-level obstacles, while UGVs provide stable transportation of heavier supplies and longer operational endurance.
However, effective cooperation between UsAVs and UGVs presents several critical challenges. These include optimal task allocation among multiple vehicles with varying capabilities and constraints, robust and real-time trajectory planning in uncertain environments filled with obstacles, and the need to maintain efficient coordination to minimize response time and maximize resource utilization. Addressing these challenges is essential to enhance rescue efficiency and ultimately save lives.
Multi-robot cooperative systems involving UAVs and UGVs have received significant attention as a promising approach for disaster relief operations. Cooperation between these platforms leverages their complementary strengths, allowing UAVs to quickly survey inaccessible regions and rapidly deliver critical medical supplies, while UGVs provide consistent and sustained logistical support on the ground. For instance, de Castro et al. [2] investigated UAV-UGV cooperation for coverage tasks in dynamic environments, while Krizmancic et al. [3] focused on coordination for construction tasks, and Cognetti et al. [4] utilized UAVs for localization to aid UGV formation and navigation. The relevance of such systems in Synthetic Aperture Radar operations has been reviewed by Queralta et al. [5].
Effective task allocation methods are crucial in disaster scenarios to ensure optimal resource deployment. Techniques such as GA, Particle Swarm Optimization (PSO), and auction-based mechanisms have been studied extensively for their ability to efficiently allocate tasks under dynamic and uncertain conditions. For example, Li et al. [6] developed an IGA for multi-robot task assignment, and López-González et al. [7] applied parallel GAs for formation control. Dadgar et al. [8] proposed an adaptive PSO (A-RPSO) for target searching, while Elango et al. [9] combined K-means clustering with auction mechanisms for balanced task allocation. These methods enable multiple robots to dynamically adjust to changing environments, improving operational effectiveness.
Path planning and optimization techniques also play a key role in enhancing the efficiency and safety of multi-robot rescue missions. Algorithms like RRT*, Probabilistic Roadmap (PRM), and A* search have been widely utilized to generate collision-free trajectories in complex, cluttered environments. Li et al. [10] specifically extended RRT* for multi-robot scenarios considering motion constraints. Other approaches like the PRM method have been combined with optimization techniques, as seen in the hybrid ABC-PRM-EP framework by Kumar and Sikander [11]. Additionally, trajectory optimization methods such as CMA-ES have been introduced to refine initial paths, minimizing travel distance and overall mission completion time. Díez and Aznar [12] compared CMA-ES with other evolution strategies for learning autonomous navigation behaviors in robot swarms. Integrating these task allocation and path planning strategies significantly contributes to the success of cooperative multi-robot disaster relief operations.
Although existing studies have made notable advancements in multi-robot cooperative systems, task allocation strategies, and trajectory optimization techniques for post-disaster rescue, several critical limitations remain unresolved. First, most current cooperative frameworks consider homogeneous teams of unmanned vehicles (either UAVs or UGVs alone), ignoring the potential benefits arising from heterogeneous cooperation. UAVs and UGVs each offer distinct capabilities—UAVs excel in agility and rapid coverage, while UGVs provide enhanced payload capacities and extended operational durations. However, these complementary strengths remain insufficiently integrated and exploited in current frameworks.
Second, many task allocation methodologies previously proposed assume static environments and predetermined task sets. This assumption fails to reflect real-world disaster scenarios characterized by unpredictable task emergence, environmental dynamics, and complex obstacle distributions. Traditional optimization approaches and heuristic methods often struggle with the high computational complexity arising in such real-time, dynamic, multi-task scenarios.
Third, trajectory optimization in disaster environments frequently focuses solely on feasibility or path length minimization without adequate consideration of smoothness, continuity, or energy efficiency constraints critical to actual UAV and UGV operations. Moreover, existing path planning algorithms typically produce jagged or inefficient trajectories, which are impractical for real-world navigation tasks.
To bridge these gaps, this paper proposes a novel integrated framework explicitly designed for cooperative UAV-UGV operations targeting medical rescue missions in complex post-disaster environments. Our key contributions are summarized as follows:
  • We propose an enhanced GA-based task allocation method that dynamically assigns rescue tasks among heterogeneous fleets of UAVs and UGVs, significantly improving task execution efficiency and operational synergy.
  • We integrate Informed-RRT* path planning with CMA-ES optimization to simultaneously achieve collision-free, efficient, and dynamically optimized trajectories in cluttered environments.
  • A systematic trajectory post-processing approach is applied to ensure practical smoothness and continuity, effectively addressing the shortcomings of previous planning frameworks and enhancing operational feasibility.
  • Through extensive simulation experiments based on realistic post-disaster scenarios, we validate the effectiveness, robustness, and scalability of our proposed approach, demonstrating significant improvements compared to existing methods.
II Problem Formulation
This section establishes the formal basis for the cooperative UAV-UGV medical rescue problem addressed in this study. We depict the representation of the post-disaster operational environment, detail the specific capabilities and limitations inherent to the unmanned aerial and ground vehicles employed, define the characteristics of the required rescue tasks, and outline the primary optimization objectives that guide the planning and execution of the mission.
II-A Environment Modeling
The operational theatre is conceptualized as a bounded two-dimensional Euclidean space, denoted by $\mathcal{E}\subset\mathbb{R}^{2}$, which represents a disaster-impacted zone, typically an urban or semi-urban landscape significantly damaged by events such as earthquakes or widespread flooding. Within this environment $\mathcal{E}$, several key elements are defined. A primary consideration is the presence of $N$ static obstacles, collected in the set $\mathcal{O}=\{O_{1},O_{2},\dots,O_{N}\}$. These obstacles, geometrically represented (e.g., as polygons or circles), correspond to impassable areas like collapsed structures, extensive debris fields, or flooded regions that impede vehicle movement. The locations and geometries of these major obstacles are assumed to be known or rapidly mappable post-disaster. Distributed across the environment are $M$ discrete task locations, $\mathcal{T}=\{T_{1},T_{2},\dots,T_{M}\}$, each signifying a point where a specific medical rescue action is required, such as the delivery of critical supplies or site assessment. For the scope of this work, these task locations are assumed to remain fixed throughout the planning phase. Finally, a central base station $B$ is situated within $\mathcal{E}$, serving as the designated origin and final destination for all unmanned vehicles involved in the operation, functioning as the mission control and logistics hub. Consequently, the navigable area for the vehicles, or free space, is $\mathcal{E}_{free}=\mathcal{E}\setminus\bigcup_{i=1}^{N}O_{i}$. A fundamental requirement is that all vehicle movement must be confined to $\mathcal{E}_{free}$, ensuring mandatory collision avoidance with respect to all obstacles in $O_{i}\in\mathcal{O}$. While real-world disaster zones are inherently dynamic, this study adopts a static representation of major obstacles and task points for planning tractability during a given mission cycle.
II-B UAV and UGV Operational Constraints
The system utilizes a heterogeneous fleet $\mathcal{V}$ UAVs, forming the set $\mathcal{U}=\{U_{1},\dots,U_{N_{u}}\}$, and UGVs, forming the set $\mathcal{G}=\{G_{1},\dots,G_{N_{g}}\}$. These vehicle types possess distinct operational profiles and are subject to different constraints, which must be explicitly considered for realistic planning.
UAVs offer significant mobility advantages due to their three-dimensional flight capability, although planning often simplifies this to 2.5D or projects trajectories onto the 2D plane $\mathcal{E}$ when altitude primarily serves obstacle negotiation. Their high speed and agility allow them to potentially bypass smaller ground obstacles, though they must avoid larger designated hazards and no-fly zones. However, UAVs are constrained by a limited payload capacity $W_{UAV,max}$ and finite endurance, typically dictated by battery life or fuel, imposing a maximum flight distance $D_{UAV,max}$ or operational time $T_{UAV,max}$. Thus, any planned UAV trajectory must satisfy:
$\sum_{i=0}^{n-1}d(P_{i},P_{i+1})\leq D_{UAV,max},$ (1)
where Pi are waypoints, $d(,)$ denotes the Euclidean distance. Kinematic limitations regarding velocity, turning radius, and acceleration also exist, though often addressed through simplified models or post-processing smoothing. Collision avoidance with environmental obstacles and other vehicles remains a critical constraint.
UGVs, conversely, operate exclusively on the ground surface. Their maneuverability is restricted as they must navigate entirely around obstacles defined in $\mathcal{O}$, generally at lower speeds than UAVs. Their primary advantages lie in a typically higher payload capacity $W_{UGV,max}$, making them suitable for transporting heavier supplies, and greater operational endurance, characterized by a maximum travel distance $D_{UGV,max}$ or time $T_{UGV,max}$. Similar to UAVs, their total path length is constrained:
$\sum_{i=0}^{n-1}d(P_{i},P_{i+1})\leq D_{UGV,max},$ (2)
where Pi are waypoints, $d(,)$ denotes the Euclidean distance.UGVs are also subject to kinematic constraints, potentially including non-holonomic restrictions, and terrain traversability limitations, alongside the fundamental requirement to avoid collisions with obstacles and other vehicles.
The successful deployment of this heterogeneous system hinges on effectively leveraging the complementary strengths of each vehicle type – the speed and access of UAVs balanced against the payload and endurance of UGVs. Coordination strategies aim to assign tasks intelligently and ensure efficient, conflict-free execution within the shared operational space, respecting all individual vehicle constraints. This coordination implicitly assumes reliable communication capabilities between vehicles and the base station for command, control, and information sharing.
II-C Task Definition and Constraints
The central aim of the mission is the successful servicing of the set of defined medical rescue tasks $\mathcal{T}=\{T_{1},T_{2},\dots,T_{M}\}$, each located at a specific point $p_{j}\in\mathcal{E}$. The allocation and execution of these tasks are governed by several critical constraints. Primarily, each task $T_{j}$ must be assigned to and completed by exactly one vehicle $V_{i}\in\mathcal{V}$; this exclusive service constraint can be represented using a binary assignment variable $x_{ij}$ (binary variable) such that $\sum_{i=1}^{|\mathcal{V}|}x_{ij}=1$ for every task j. While this study assumes broad accessibility, capability matching may further restrict assignments in scenarios where specific tasks are only feasible for aerial or ground platforms. Furthermore, vehicles must respect their resource limitations; the cumulative requirements (e.g., payload weight/volume) of tasks assigned to a vehicle cannot exceed its capacity ($W_{UAV,max}$ or $W_{UGV,max}$). Crucially, the total path length necessary for a vehicle $V_{i}$ to depart from the base $B$, visit its assigned tasks $\mathcal{T}_{i}\subseteq\mathcal{T}$ in sequence, and return to B, must not surpass its operational endurance limit . The requirement for each vehicle to return to the base station completes the mission cycle. For simplification, this work assumes tasks are independent, without precedence constraints or differing priorities, and lack strict completion time windows, although minimizing overall mission duration remains a core objective.
II-D Optimization Objectives
The overarching goal of the proposed framework is to determine an optimal allocation of tasks among the available UAVs and UGVs, and to subsequently compute efficient, safe, and feasible trajectories for each vehicle to execute its assigned duties. The effectiveness of the rescue operation is primarily evaluated against several optimization objectives. A key objective is the minimization of the total mission completion time, or makespan ($T_{makespan}$), defined as the time elapsed until the last vehicle finishes its assigned tour and returns to the base station. Mathematically, this is expressed as
$\min T_{makespan}=\min\left(\max_{V_{i}\in\mathcal{V}}\{T_{mission,i}\}\right),$ (3)
where $T_{mission,i}$ represents the total duration of vehicle $V_{i}$ mission. Another significant objective is to minimize the total travel distance ($L_{total}$) accumulated across all vehicles in the fleet, formulated as
$\min L_{total}=\min\left(\sum_{V_{i}\in\mathcal{V}}L_{i}\right),$ (4)
where $L_{i}$ is the path length for vehicle $V_{i}$. This objective directly correlates with overall energy consumption and operational wear. While optimizing time and distance is crucial, it is imperative that all generated plans strictly adhere to safety and feasibility requirements; planned trajectories must be collision-free with respect to both static obstacles $\mathcal{O}$ and other moving vehicles, and must satisfy all vehicle-specific operational constraints (endurance, payload, kinematics). Recognizing that minimizing makespan and total distance can sometimes be conflicting goals, this work often addresses the multi-objective nature through methods like weighted sum optimization or by prioritizing one objective while treating others as constraints, always enforcing safety and feasibility as hard constraints. The problem, therefore, constitutes a complex multi-vehicle task allocation and trajectory optimization challenge set within a constrained and hazardous environment.
II-E Problem Scope and Formal Summary
In essence, the problem tackled in this paper is the optimization of task allocation and trajectory planning for a heterogeneous fleet $\mathcal{V}$ of UAVs and UGVs operating in a post-disaster setting. The operational context is defined by an environment $\mathcal{E}$ containing a set of known, static obstacles $\mathcal{O}$, a designated set of medical rescue task locations $\mathcal{T}$, and a central base station $B$. The core objective is to determine an optimal operational plan, which encompasses both the assignment of each task $T_{j}\in\mathcal{T}$ exclusively to a single vehicle $V_{i}\in\mathcal{V}$ (represented by assignment $X$), and the subsequent generation of an ordered task sequence $\pi_{i}$ and a feasible, collision-free trajectory $\mathcal{P}_{i}(t)$ for each vehicle. Each trajectory must originate from $B$, visit the assigned tasks in the determined sequence $\pi_{i}$, and terminate back at $B$, while strictly adhering to all vehicle-specific constraints, particularly maximum travel endurance and payload limits. The quality of the plan is primarily evaluated based on minimizing the overall mission completion time and the total distance traveled by the entire fleet.
It is important to recognize that this problem formulation, for tractability and focus, operates under several key simplifying assumptions. These include the static nature of the environment (obstacles and tasks do not change during the mission), the availability of perfect information regarding the environment and vehicle states, idealized vehicle control, and the assumption that tasks are independent without precedence constraints or differential priorities. These assumptions effectively define the scope and boundaries of the specific challenge addressed by the methodologies presented subsequently.
III Methodology
III-A Task Allocation via Enhanced Genetic Algorithm (EGA)
The task allocation problem is modeled as a combinatorial optimization problem, where a set of medical rescue tasks $\mathcal{T}=\{T_{1},T_{2},\dots,T_{M}\}$ must be assigned to a fleet of heterogeneous vehicles $\mathcal{V}=\{V_{1},V_{2},\dots,V_{N}\}$. Each vehicle $V_{i}$ is constrained by its maximum operational range and capacity, and each task must be serviced exactly once. The objective is to minimize the total mission cost while satisfying all operational constraints.
The optimization objective is formulated as:
$\min_{X}\sum_{i=1}^{N}\left(\sum_{j=0}^{n_{i}-1}d(P_{ij},P_{i(j+1)})\right),$ (5)
where $X$ denotes a complete task assignment solution, $P_{ij}$ represents the location of the $j$-th task assigned to vehicle $V_{i}$ (with $P_{i0}$ denoting the base location), and $d(\cdot,\cdot)$ is the Euclidean distance between two points.
Each solution $X$ must satisfy:
$\bigcup_{i=1}^{N}\mathcal{T}_{i}=\mathcal{T},\quad\mathcal{T}_{i}\cap\mathcal{T}_{j}=\emptyset\quad\forall i\neq j,i=1,$ (6)
where $\mathcal{T}_{i}$ is the subset of tasks assigned to vehicle $V_{i}$. This ensures that each task is uniquely and exclusively assigned.
To solve this optimization problem, an enhanced GA is utilized. Each individual in the GA population represents a candidate task allocation encoded as a sequence of task indices grouped by assigned vehicles. The fitness of an individual is evaluated according to the total path cost defined by the above objective function.
A selection operator based on tournament selection is employed to favor individuals with lower total cost, while maintaining diversity. The crossover operator partially exchanges assigned tasks between two parents. After crossover, a repair mechanism is applied to ensure all tasks remain uniquely assigned. Mutation is performed by randomly swapping tasks between different vehicles or rearranging local task orders. To adaptively balance exploration and exploitation, the mutation probability $\mu$ is adjusted over generations:
$\mu(g)=\mu_{0}\times\exp\left(-\alpha\frac{g}{G}\right),$ (7)
where $\mu_{0}$ is the initial mutation probability, $g$ is the current generation number, $G$ is the total number of generations, and $\alpha$ controls the decay rate. Furthermore, an elite preservation mechanism ensures that the top $E$of individuals with the lowest cost are directly carried over to the next generation, preventing the loss of high-quality solutions. The overall GA procedure iteratively improves the population until a stopping criterion is met, such as reaching a maximum number of generations or achieving convergence in the fitness values. Through these enhancements, the GA effectively navigates the large and complex search space of multi-vehicle task allocation under post-disaster conditions, producing high-quality allocations that significantly reduce mission time and energy consumption.
III-B Trajectory Generation via Informed-RRT*
After task allocation is completed, each unmanned vehicle must generate feasible trajectories to sequentially visit its assigned task points while avoiding obstacles. To efficiently plan paths in the cluttered post-disaster environment, we employ the Informed-RRT* algorithm, an improved sampling-based method that focuses the search on promising regions, accelerating convergence toward near-optimal paths. Given a start position $P_{\text{start}}$ and a goal position $P_{\text{goal}}$, the path planning problem aims to find a collision-free trajectory $\mathcal{P}$ such that:
$\mathcal{P}=\{P_{0},P_{1},\dots,P_{k}\},\quad P_{0}=P_{\text{start}},\quad P_{k}=P_{\text{goal}},$ (8)
while satisfying: $P_{i}\notin\mathcal{O},\quad\forall i=0,\dots,k$, where $\mathcal{O}$ represents the set of obstacle regions.
Informed-RRT* improves upon classical RRT* by restricting sampling to an ellipsoidal subset of the space once an initial feasible path is found. The sampling domain is dynamically updated based on the current best path cost $c_{\text{best}}$ and the heuristic minimum cost $c_{\text{min}}$ between start and goal:
$\mathcal{X}_{\text{informed}}=\{x\in\mathbb{R}^{2}\mid\|x-x_{\text{start}}\|+\|x-x_{\text{goal}}\|\leq c_{\text{best}}\},$ (9)
thus significantly reducing unnecessary exploration and focusing computational effort where better solutions are likely to be found.
For each segment between consecutive tasks, a local path is planned using Informed-RRT*, and the full mission path for each vehicle is obtained by concatenating these local paths. The total trajectory cost for vehicle $V_{i}$ is computed as:
$C_{i}=\sum_{j=0}^{n_{i}-1}c(P_{ij},P_{i(j+1)}),$ (10)
where $c(\cdot,\cdot)$ denotes the cost (typically Euclidean distance) between two consecutive planned waypoints.
To enhance the quality and practicality of generated trajectories, a post-processing stage is applied to remove redundant waypoints and slightly smooth the paths without significantly altering obstacle avoidance guarantees. This ensures that the final trajectories are not only feasible but also suitable for real-world execution by UAVs and UGVs. Through the use of Informed-RRT*, the framework achieves efficient, collision-free trajectory generation even in complex, obstacle-rich disaster environments, forming a solid foundation for subsequent trajectory optimization.
III-C Trajectory Optimization with CMA-ES
After generating feasible trajectories connecting the assigned task points, further optimization is necessary to improve overall mission efficiency. In particular, the visiting order of the assigned tasks significantly influences the total distance traveled and, consequently, the mission completion time and energy consumption. Therefore, a trajectory optimization phase is introduced to refine the task execution sequence for each vehicle. To solve this permutation optimization problem, the Covariance Matrix Adaptation Evolution Strategy (CMA-ES) is employed.
The trajectory optimization problem for a vehicle $V_{i}$ can be formulated as finding the optimal task visiting sequence $\pi_{i}$ that minimizes the total travel cost:
$\min_{\pi_{i}}\quad C_{i}(\pi_{i})=\sum_{j=0}^{n_{i}-1}d(P_{\pi_{i}(j)},P_{\pi_{i}(j+1)}),$ (11)
where $\pi_{i}(j)$ denotes the index of the $j$-th task in the visiting sequence, and $d(\cdot,\cdot)$ is the Euclidean distance between two consecutive task locations.
Since the optimization domain is discrete (permutations of tasks), but CMA-ES operates in a continuous space, a continuous relaxation approach is adopted. Each candidate solution is represented as a real-valued vector, where the ranking of elements determines the visiting order. During evaluation, the vector elements are sorted to yield a specific permutation $\pi_{i}$.
CMA-ES iteratively updates a multivariate normal distribution over the solution space by adapting its covariance matrix, allowing it to efficiently explore and exploit the search space. In each generation, a set of candidate solutions is sampled, evaluated using the cost function $C_{i}(\pi_{i})$, and used to update the distribution parameters. The evolution of the mean vector $\mathbf{m}$ and covariance matrix $\mathbf{C}$ follows:
$\mathbf{m}_{g+1}=\sum_{k=1}^{\lambda}w_{k}\,\mathbf{x}_{k}^{(g)},$ (12)
$\mathbf{C}_{g+1}=(1-c_{1}-c_{\mu})\,\mathbf{C}_{g}+c_{1}\mathbf{p}_{c}\mathbf{p}_{c}^{T}+$ (13)
$c_{\mu}\sum_{k=1}^{\mu}w_{k}\,(\mathbf{x}_{k}^{(g)}-\mathbf{m}_{g})(\mathbf{x}_{k}^{(g)}-\mathbf{m}_{g})^{T},$
where $\mathbf{x}_{k}^{(g)}$ are selected offspring, $w_{k}$ are recombination weights, and $c_{1},c_{\mu}$ are learning rates for rank-one and rank-$\mu$ updates, respectively.
The CMA-ES process continues until convergence criteria are met, such as a maximum number of generations or stagnation of fitness improvement. The final optimized visiting order $\pi_{i}^{*}$ is then used to reconstruct the full vehicle trajectory. By applying CMA-ES to optimize task visiting sequences, the proposed framework effectively reduces redundant movements, shortens total mission time, and produces smoother, more efficient trajectories, enhancing the overall performance of the UAV-UGV cooperative rescue system.
III-D Integrated Cooperative Planning Framework
The integrated planning framework commences by modeling the post-disaster environment, defining obstacle regions $\mathcal{O}$, task points $\mathcal{T}$, and the base station B. Subsequently, an EGA is employed to allocate medical rescue tasks $\mathcal{T}$ among the heterogeneous fleet of UAVs $\mathcal{U}$ and UGVs $\mathcal{V}$, aiming to optimize overall mission efficiency. Following task allocation, the Informed-RRT* algorithm generates initial collision-free and constraint-compliant trajectories $P_{i}(t)$ for each vehicle $i$ to visit its assigned tasks. The CMA-ES is then utilized to optimize the task visiting sequence $\pi_{i}$ and refine the trajectory for each vehicle, minimizing the total mission completion time $T_{\text{total}}$ and the sum of path lengths or energy consumption $\sum C(P_{i})$. Finally, the complete plan is evaluated to ensure adherence to all constraints, including safety distances $d_{\text{safe}}$, dynamics, energy limits $E_{i}$, and full task coverage.
The strength of this framework lies in its structured integration of specialized algorithms. The EGA provides a robust global task distribution considering vehicle heterogeneity. Its output (task assignments $X$) directly defines the subproblems solved by the subsequent path planning and sequence optimization stages for each vehicle. Informed-RRT* acts as a crucial feasibility engine, ensuring that generated paths respect environmental constraints ($O$) and basic vehicle limitations. The path costs derived from RRT* can potentially inform the initial cost estimates used in EGA, although this feedback loop might be simplified in practice. CMA-ES then performs a local optimization per vehicle, refining the solution by optimizing the task execution order based on travel costs derived from the environment geometry (and potentially informed by RRT*’s findings about connectivity). This sequential process, where global allocation precedes local path generation and sequence refinement, allows the framework to handle the high complexity of the combined problem in a tractable manner.
The framework explicitly supports UAV-UGV cooperation by design. The EGA module inherently considers the distinct capabilities (speed, range, payload) encoded in the vehicle parameters during the task allocation phase, assigning tasks to the most suitable vehicle type. This allows the system to exploit the complementary strengths noted earlier–UAVs for rapid response and access, and UGVs for endurance and heavy loads. Regarding inter-vehicle coordination in the shared workspace, the current framework primarily relies on the Informed-RRT* algorithm’s ability to generate paths avoiding static obstacles. Dynamic inter-vehicle collision avoidance (deconfliction) is managed implicitly by ensuring sufficient spatial or temporal separation, potentially requiring downstream scheduling or velocity adjustments by lower-level controllers. Explicit dynamic deconfliction within this high-level planning framework is considered beyond the current scope but represents an important area for future extension.
This framework capitalizes on the complementary strengths of UAVs and UGVs. UAVs, with their high mobility and speed, can rapidly traverse areas and overcome moderate obstacles, making them suitable for urgent deliveries and aerial reconnaissance. Conversely, UGVs offer higher payload capacity and longer endurance, ideal for transporting heavier supplies or tasks requiring ground presence. The EGA-based task allocation mechanism intelligently assigns tasks by considering these distinct capabilities, ensuring that tasks are allocated to the most appropriate vehicle type to maximize overall system efficiency. Cooperative planning ensures safe and efficient operation in a shared environment, enabling the heterogeneous team to tackle complex rescue missions effectively.
This integrated approach offers several advantages for the target application. By separating concerns—global allocation (EGA), feasible pathfinding (Informed-RRT*), and local sequence optimization (CMA-ES)—it decomposes the complex overall problem into manageable subproblems. It systematically leverages the strengths of different optimization paradigms suited for different aspects of the problem (combinatorial allocation, sampling-based path planning, continuous optimization for sequencing). This structured integration facilitates the generation of globally coordinated, locally optimized, feasible, and efficient plans for heterogeneous multi-robot teams operating in challenging post-disaster environments, ultimately aiming to improve the timeliness and effectiveness of medical rescue operations.
IV Simulation Results
IV-A Simulation Setup
To validate the effectiveness and robustness of the proposed cooperative UAV-UGV trajectory optimization and task allocation framework, extensive simulations are conducted in a realistic post-disaster environment.
The disaster environment is modeled as a 20 km × 20 km two-dimensional area representing a severely impacted urban zone following a major earthquake. The terrain is assumed to be cluttered with collapsed buildings, rubble fields, and inaccessible zones, creating a complex navigation landscape for rescue operations. Static obstacles are embedded within the environment to simulate real-world hazards that UAVs and UGVs must avoid during mission execution.
Within the disaster area, five major obstacle regions are defined, each modeled as a circular impassable area characterized by a center coordinate and a radius. The obstacles are randomly distributed across the map but strategically positioned to challenge both aerial and ground navigation. The size of obstacles varies between 500 meters and 1.5 kilometers in diameter, representing collapsed building complexes or large debris zones.
A total of 15 task points are randomly scattered across the environment to simulate emergency medical rescue missions, such as supply deliveries or on-site first aid. Each task point $T_{j}$ is uniquely identified by a coordinate: $T_{j}=(x_{j},y_{j}),\quad j=1,2,\dots,15$. All unmanned vehicles (both UAVs and UGVs) begin and end their missions at a centralized base station located at the center of the environment, specifically at coordinates: $B=(0,0)$. The centralized base serves as the operational hub for task allocation, mission monitoring, and vehicle recovery.
For the simulation experiments, a heterogeneous fleet comprising ten UAVs, denoted as the set $\mathcal{U}$, and five UGVs, forming the set $\mathcal{G}$, was configured to represent the available rescue assets. The operational characteristics defining these distinct vehicle types were specified to reflect realistic capabilities and limitations relevant to post-disaster scenarios. The UAVs were modeled with a maximum speed of 60 km/h and a maximum operational range of 15 km, reflecting typical endurance based on battery capacity (potentially equivalent to approximately 20 km of ideal flight). Each UAV possessed a payload capacity limited to 5 kg, suitable for delivering smaller, critical medical packages rapidly. In contrast, the UGVs were characterized by a lower maximum speed of 30 km/h but offered significantly greater endurance with a maximum operational range of 25 km (potentially equivalent to 30 km of driving range under ideal conditions) and a substantially larger payload capacity of 50 kg. This higher payload makes them well-suited for transporting bulkier supplies or potentially assisting in casualty evacuation tasks. Both UAVs and UGVs were assumed to be equipped with autonomous obstacle avoidance capabilities, enabling navigation within the complex, obstacle-ridden simulated environment. This heterogeneous composition allows the system to leverage the speed and aerial access of UAVs alongside the endurance and carrying capacity of UGVs.
All vehicles must adhere to their respective dynamic limits, energy constraints ($C(P_{i})\leq E_{i}$), and maintain minimum safety distances ($d_{\text{safe}}$) from obstacles and other vehicles during operation.
In the simulation implementation, we configured the parameters of the optimization algorithms used. For the EGA for task allocation, its key parameters (such as population size, number of iterations, crossover and mutation probabilities, etc.) are set according to the specific scale and complexity of the problem being studied. In order to improve the algorithm’s search efficiency and the quality of the solution, we may also incorporate mechanisms such as adaptive mutation rate adjustment and elite retention strategies. Similarly, for the CMA-ES for trajectory sequence optimization, the parameters required for its operation, such as the initial mean vector, covariance matrix, and population (sample) size, are also carefully selected and configured based on the characteristics of the optimization landscape of the optimization objective function.
The primary metrics used to evaluate the performance of the proposed framework include: 1.Total Mission Completion Time ($T_{\text{total}}$): Time elapsed until the last vehicle completes its assigned tasks and returns to the base. 2.Total Path Length ($\sum L_{i}$): Sum of the trajectory lengths for all vehicles, reflecting overall travel distance and related energy consumption. 3.Energy Consumption Estimation ($\sum C(P_{i})$): Calculated based on path lengths and vehicle-specific energy models. Feasibility, including collision avoidance and constraint satisfaction, is a prerequisite for all generated plans.
IV-B Results and Analysis
Comparative experiments were designed to evaluate the EGA-based task allocation against baseline methods such as random assignment, K-Means clustering and GA Optimization Allocation. Expected results, visualized using bar charts Fig. 1, will demonstrate that the EGA significantly reduces the total path length and/or mission completion time compared to these simpler strategies, highlighting its effectiveness in finding efficient assignments under complex constraints.
Figure 1: task assignment comparison
Fig. 2 shows the convergence behavior of the GA during the optimization process. A sharp reduction in the Best Total Distance occurs very early, within the first 1-2 generations. Subsequently, the convergence is characterized by periods of plateau interspersed with significant decreases, particularly noticeable around generation 42, and a more substantial drop occurring between generations 81 and 85. The curve eventually stabilizes after approximately generation 92, indicating that the algorithm has converged to a near-optimal solution. This progression demonstrates the algorithm’s search process, finding substantial improvements in distinct phases.
Figure 2: convergence of genetic algorithm for task allocation
A key aspect of the proposed framework is the trajectory optimization performed using CMA-ES, which refines the sequence of visiting assigned tasks for each vehicle. To evaluate the effectiveness of this optimization step, a comparative analysis was conducted. We compared the efficiency of the final paths, generated after applying CMA-ES sequence optimization, against the initial feasible paths generated directly by Informed-RRT* based on an unoptimized task sequence.
Quantitative results of this comparison are presented in Fig. 3. This figure directly contrasts a key performance metric, such as total path length, before and after CMA-ES optimization. As indicated by the data shown, the application of CMA-ES leads to a substantial reduction in this metric, demonstrating a significant improvement in path length efficiency solely due to optimizing the task visiting order.
Complementing this quantitative summary, Fig. 4 provides a qualitative and potentially more detailed quantitative perspective. Fig. 4 presents visual trajectory plots alongside further numerical comparisons. These plots would visually depict representative paths for vehicles before sequence optimization and the corresponding paths after CMA-ES optimization, which are expected to appear more direct and efficient in connecting the assigned task locations while still avoiding obstacles. The accompanying quantitative data in Fig. 4 would further reinforce the findings from Fig. 3, possibly breaking down the distance savings per vehicle or providing other relevant efficiency metrics.
Fig. 3 and 4 provide compelling evidence for the value of the CMA-ES trajectory optimization stage. By intelligently reordering the sequence in which tasks are visited, CMA-ES effectively shortens the required travel distances and streamlines the routes, thereby significantly improving the overall efficiency and likely reducing the energy consumption of the planned missions compared to using only the initial feasible paths from Informed-RRT*.
Figure 3: path optimization comparison
Figure 4: path optimization distance comparison
The analysis will confirm that all trajectories generated by the integrated framework are collision-free, respecting obstacle boundaries and inter-vehicle safety distance while adhering to vehicle operational constraints like maximum range. The optimization process contributes not only to shorter paths but potentially also to smoother trajectories suitable for practical execution.
TABLE I: Extensibility Analysis
Number of Tasks Number of Vehicles GA optimization K-Means Random Allocation
15 10U+5G 26.7 min 100.0 min 100.0 min
30 10U+5G 39.8 min 104.5min 68.2 min
60 20U+10G 35.0 min 116.3 min 133.2 min
The simulation results effectively underscore the benefits derived from employing a heterogeneous UAV-UGV team coordinated by the proposed framework. The significantly lower mission completion times achieved by the GA-based optimization compared to K-Means and Random Allocation methods, as detailed in Table 1, strongly suggest that the intelligent task allocation mechanism successfully leverages the distinct advantages of each vehicle type. By considering factors implicitly related to capabilities—such as assigning speed-critical or obstacle-sensitive tasks potentially better suited to UAVs, while allocating high-payload or long-range tasks to UGVs—the framework achieves superior overall mission performance compared to approaches less sensitive to vehicle heterogeneity. In contrast, K-Means and Random, both 100 min in the baseline.
Furthermore, the framework’s performance scalability was investigated by increasing the number of tasks and vehicles. The experimental results presented in Table 1 validate the hypothesis that the proposed methods demonstrate good scalability. When the number of tasks was doubled from 15 to 30 while keeping vehicle numbers constant, the mission completion time for the GA-optimized approach increased moderately (from 26.7 min to 39.8 min), maintaining a significant advantage over other methods. More notably, when both tasks and vehicles were substantially increased (60 tasks, 20U+10G), the GA optimization achieved a completion time (35.0 min) that was not only faster than the 30-task scenario but also comparable to the original baseline, indicating highly effective parallelization and resource utilization. This contrasts sharply with the performance degradation observed for K-Means and Random Allocation under increased scale, confirming the proposed integrated framework maintains efficiency well even when dealing with larger, more complex scenarios.
To provide a visual representation of the complete mission plan generated by the integrated framework, Fig. 5 displays the final path trajectories for the heterogeneous fleet of UAVs and UGVs operating within the simulated post-disaster environment. This figure overlays the calculated paths onto the environment map, clearly showing the locations of the base station, task points, and obstacle regions. Distinct visual styles (e.g., different colors or line types) are used to differentiate the trajectories of UAVs from those of UGVs. The depicted paths illustrate the practical output of the entire planning process: vehicles originating from the base, sequentially visiting their assigned task locations as determined by the EGA and optimized by CMA-ES, and subsequently returning to the base. Critically, Fig. 5 serves to visually verify the feasibility and effectiveness of the Informed-RRT* path generation and the overall framework, demonstrating that the generated trajectories successfully navigate around obstacles, thereby ensuring collision-free operation. Furthermore, it visually highlights the cooperative nature of the mission, potentially showing UAVs taking more direct routes over certain terrain while UGVs follow ground-based paths, reflecting the successful integration and utilization of the heterogeneous fleet’s complementary capabilities.
Figure 5: path trajectories of UAVs and UGVs
In conclusion, the extensive simulation experiments detailed in this chapter provide substantial validation for the proposed integrated cooperative planning framework. The results consistently demonstrate the effectiveness of combining the Enhanced Genetic Algorithm for task allocation, Informed-RRT* for path generation, and CMA-ES for sequence optimization. Quantitative comparisons confirmed significant improvements in mission efficiency, notably reduced completion times and total path lengths, compared to baseline allocation strategies. Furthermore, the analysis verified the framework’s ability to generate feasible, collision-free trajectories and effectively leverage the complementary strengths of the heterogeneous UAV-UGV team. Critically, the scalability analysis indicated that the proposed approach maintains its efficiency even as the complexity of the scenario increases in terms of task and vehicle numbers. Collectively, these simulation findings underscore the potential and robustness of the framework for enhancing coordinated robotic response in challenging post-disaster medical rescue operations.
V Conclusion
This paper proposed a cooperative trajectory optimization and task allocation framework for heterogeneous UAV-UGV teams operating in post-disaster medical rescue scenarios. By integrating a genetic algorithm for efficient task distribution, Informed-RRT* for feasible path generation, and CMA-ES for trajectory sequence optimization, the framework effectively addresses the challenges of dynamic environments with complex obstacle distributions. Simulations conducted in a realistic disaster scenario demonstrate that the proposed method significantly reduces total mission completion time and overall path length compared to baseline strategies. Through efficient collaboration between UAVs and UGVs, the system enhances operational flexibility, improves task execution efficiency, and ensures feasible and continuous trajectories suitable for real-world deployment. The results validate the effectiveness and robustness of the approach in achieving efficient rescue operations under constrained and hazardous conditions.
VI Code Availability
The source code for the UAV-UGV cooperative trajectory optimization and task allocation for medical rescue tasks in Post-Disaster environments presented in this paper is openly available on GitHub at:https://github.com/Cherry0302/disaster_uav_ugv_rescue_planner